OpenStudy (anonymous):

If sin y°=and tan y°=,what is the value of cos y°? cos y° =6/b cos y°= 6a cos y°= 6b cos y° =b/6 i do not get this soh cah toah or whatever

OpenStudy (dumbcow):

SOH CAH TOA my fav acronym in math :) i think you are missing info sin y = ?

OpenStudy (anonymous):

sorry it didn't load. If sin y°=a/6and tan y°=a/b,what is the value of cos y°?

OpenStudy (dumbcow):

for this question, just use the fact that tan = sin/cos --> a/b = (a/6)/cos --> cos = (a/6)/(a/b) = a/6 * b/a = b/6

OpenStudy (dumbcow):

 sin y = opp/hyp = a/6 cos y = adj/hyp = b/6 tan y = opp/adj = a/b